ok. i need you to help me solve 1 and 3/4 + ? = 3lbs.
Answer by edjones(8007)

1 3/4 + x = 3 assuming this to be the question.
7/4 + x =3
x=3- 7/4
x= 12/4 - 7/4
x=5/4
.
Ed